OpenStudy (crashonce):

x^2+y^2 = k^2+1 and xy=k then x-y=?

OpenStudy (crashonce):

@mathmath333 help pls I have more questions after

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{x^2+y^2=k^2+1\\~\\ x^2-2xy+y^2=k^2-2xy+1\\~\\ (x-y)^2=k^2-2k+1\\~\\ (x-y)^2=(k-1)^2\\~\\ }\end{align}\)